OpenStudy (anonymous):

In how many ways can the letters of the word number be arranged if the N is somewhere to the left of U?

OpenStudy (anonymous):

what is the word?

OpenStudy (anonymous):

sorry lol i thought i put it in. number

OpenStudy (anonymous):

oh wait i did lol

OpenStudy (anonymous):

oh wow i didnt see it at all lol

OpenStudy (anonymous):

ok i know its 24 x 15 but im not sure about the method i did i think theres an easier way

OpenStudy (akshay_budhkar):

where is the word?

OpenStudy (anonymous):

number

OpenStudy (anonymous):

i probably should have caps locked it lol

OpenStudy (akshay_budhkar):

oh number lol!!!

OpenStudy (anonymous):

yeah u got it:)

OpenStudy (anonymous):

First i would find out how many ways you can place the N and the U with the N to the left of U. Thats going to be: \[\left(\begin{matrix}6 \\ 2\end{matrix}\right) = 15\]

OpenStudy (anonymous):

oh wait we can use a formula it ends up being 6!/2! right

OpenStudy (anonymous):

Then i would look at the other four letters, and ask myself how many ways can i rearrange those letters. That would be 4! = 24 So 24*15 is the total number of ways you could rearrange the letters with those requirements.

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

so: \[\left(\begin{matrix}6 \\ 2\end{matrix}\right) = \frac{6!}{2!(6-2)!} = \frac{6*5}{1*2} = 15\]

OpenStudy (akshay_budhkar):

Master joe deserves to be one lol

OpenStudy (anonymous):

and another one :)

OpenStudy (anonymous):

hey akshay!

OpenStudy (akshay_budhkar):

hey joe where have you been?

OpenStudy (anonymous):

asleep when im supposed to be lolol

OpenStudy (akshay_budhkar):

oh i am missing you Sir!!!!!!!! Can you see i have reached level 47?

OpenStudy (anonymous):

nice nice! *tear* the kids, they grow so fast these days lolol jk

OpenStudy (akshay_budhkar):

lol